A person pushing a lawnmower on a level lawn applies a constant force of 250 N at an

angle θ below the horizontal. If the person’s push does work of 1.06 x 10 4 J on the mower and the mower moves a distance of 44.5 m, what is the angle between the directions of the push and the displacement?


1
Expert's answer
2022-04-26T14:55:47-0400

"A=Fs\cos\theta,"

"\theta=\arccos\frac A{Fs}=17.7°."